# OmniNav: A Unified Framework for Prospective Exploration and Visual-Language Navigation

### v1

@@ 4 DATA and Training
Data in the embodied domain is typically organized as a data pyramid (Bjorck et al., 2025), with internet data and human video data at the bottom, simulation or synthetic data in the middle, and real-robot data at the top. Our dataset follows this rich composition as well, including general web data, simulation data, and a very small amount of real-robot data detailed in Fig. 3
@@ 4.1 General Dataset
It is found in our experiment that models can learn the navigation paradigm relatively easily, whereas general-purpose capabilities remain challenging. To strengthen these abilities, we extended the training data with broad-coverage general-purpose datasets. General-purpose datasets (general QA, image captioning, OCR, chart understanding, coding, and math) complement Vision-and-Language Navigation (VLN) by supplying foundational skills in language understanding, visual semantics, OCR recognition, structured reasoning, and algorithmic planning. These capabilities improve instruction comprehension, path planning, and overall robustness. They also introduce commonsense and functional priors, such as “bath towels are commonly found in bathrooms”. We draw on the open-source MAmmoTH-VL corpus (Guo et al., 2024), and subsampled examples to expand our general capabilities; the composition ratios are shown in Fig. 3.
We further incorporated grounding and referring data to more reliably ground linguistic targets and relations to concrete pixels, instances, and locations in the scene, thereby yielding policies that are more interpretable, robust, and generalizable. This includes fine-grained language-to-target mapping—for example, precisely localizing language like “red sofa,” “door with a handle,” or “the second chair” to the correct image regions to support instance-level disambiguation. It also encompasses spatial and relational understanding, such as learning to ground spatial prepositions and relations (“to the left/in front of/at the corner”), and extracting actionable turning and stopping cues from relational descriptions like “the door next to the painting” or “after passing the hallway, turn right at the first intersection.” The referring and grounding data are sourced from RefCOCO series (Chen et al., 2025b) and Objects365 (Shao et al., 2019).
@@ 4.2 Multi-Task Navigation Dataset
Point-goal data. We use the open-source Citywalker corpus (Liu et al., 2025), which consists of first-person city-walking videos collected from public YouTube channels. Each trajectory comprises a single forward-view sequence. Long videos are segmented into 2-minute clips, then the camera pose for every frame are recovered by DPVO (Deep Patch Visual Odometry) (Teed et al., 2023).
Instruct-goal data. Under Habitat’s VLN-CE (continuous environment) setting and Matterport3D scens (Chang et al., 2017), based on the public instruction–path pairs from R2R (Anderson et al., 2018) and RxR (Ku et al., 2020). For each trajectory, we store a panoramic sequence composed of three-views (front, left, right), the action and waypoint sequences, and the natural-language instruction.
Object-goal data. Specifically, open-vocabulary object navigation (OVON) data (Yokoyama et al., 2024b). In Habitat-Matterport3D (HM3D) (Ramakrishnan et al., 2021) scenes, we randomly sample (start pose, target object category) pairs. The built-in shortest-path navigation algorithm is used to move the agent to the vicinity of the target. The same content are recorded as the instruct-goal data for each trajectory.
Object-goal (with frontier-based exploration) data. At each step the agent updates its occupancy map and visible region to identify all current frontiers. A policy then selects one frontier—favoring the shortest path while introducing limited randomness—as the current subgoal. Once the target object is found, the exploration episode terminates successfully, yielding a trajectory (Zhu et al., 2025). Each trajectory record includes a single forward-view frame sequence, a unexplored frontiers sequence and a natural-language instruction.
Embodied QA. ScanQA (Azuma et al., 2022) focuses on real-world indoor 3D scene understanding, with QA pairs centered on object locations, attributes, and spatial relations. R2R-EnvDrop (Tan et al., 2019) addresses continuous visual navigation: scenes are from Matterport3D, trajectories are from R2R-CE, and the navigation instruction–trajectory pairs are recast into a QA format to strengthen alignment between linguistic expressions and visual observations.
@@ 4.3 Discrete and Continuous Joint Training
Jointly training scheme is particularly critical for Vision-and-Language Navigation (VLN), which requires strong general knowledge and semantic understanding, leading to substantial improvements in success rates in open environments. Stage 1 is trained with 96 NVIDIA H20 GPUs for 120 hours, and Stage 2 is trained with 64 NVIDIA H20 GPUs for 48 hours with lower learning rate.
@@ 5 Experiments
HPN+DN* (Krantz et al., 2021) 6.31 40.0 36.0 34.0 - - -
CMA* (Hong et al., 2022) 6.20 52.0 41.0 36.0 8.76 26.5 22.1
Sim2Sim* (Krantz & Lee, 2022) 6.07 52.0 43.0 36.0 - - -
GridMM* (Wang et al., 2023b) 5.11 61.0 49.0 41.0 - - -
DreamWalker* (Wang et al., 2023a) 5.53 59.0 49.0 44.0 - - -
Reborn* (An et al., 2022) 5.40 57.0 50.0 46.0 5.98 48.6 42.0
ETPNav* (An et al., 2024) 4.71 65.0 57.0 49.0 5.64 54.7 44.8
HNR* (Wang et al., 2024) 4.42 67.0 61.0 51.0 5.50 56.3 46.7
AG-CMTP (Chen et al., 2021) 7.90 39.0 23.0 19.0 - - -
R2R-CMTP (Chen et al., 2021) 7.90 38.0 26.0 22.0 - - -
InstructNav (Long et al., 2024a) 6.89 - 31.0 24.0 - - -
LAW (Raychaudhuri et al., 2021) 6.83 44.0 35.0 31.0 10.90 8.0 8.0
CM2 (Georgakis et al., 2022) 7.02 41.0 34.0 27.0 - - -
WS-MGMap (Chen et al., 2022) 6.28 47.0 38.0 34.0 - - -
AO-Planner (Chen et al., 2025a) 5.55 59.0 47.0 33.0 7.06 43.3 30.5
Seq2Seq (Krantz et al., 2020) 7.77 37.0 25.0 22.0 12.10 13.9 11.9
CMA (Krantz et al., 2020) 7.37 40.0 32.0 30.0 - - -
NaVid (Zhang et al., 2024b) 5.47 49.0 37.0 35.0 - - -
Uni-NaVid (Zhang et al., 2024a) 5.58 53.5 47.0 42.7 6.24 48.7 40.9
NaVILA (Cheng et al., 2024) 5.22 62.5 54.0 49.0 6.77 49.3 44.0
StreamVLN (Wei et al., 2025) 4.98 64.2 56.9 51.9 6.22 52.9 46.0
CorrectNav (Yu et al., 2025) 4.24 67.5 65.1 62.3 4.09 69.3 63.3
@@
Method Val-Seen Val-Seen-Synonyms Val-Unseen
@@
BC (Pomerleau, 1988) 11.1 4.5 9.9 3.8 5.4 1.9
DAgger (Ross et al., 2011) 11.1 4.5 9.9 3.8 5.4 1.9
RL (Schulman et al., 2017) 18.1 9.4 15.0 7.4 10.2 4.7
DAgRL (Chen et al., 2019) 41.3 21.2 29.4 14.4 18.3 7.9
BCRL (Wang et al., 2019) 39.2 18.7 27.8 11.7 18.6 7.5
VLFM* (Yokoyama et al., 2024a) 35.2 18.6 32.4 17.3 35.2 19.6
DAgRL+OD (Yokoyama et al., 2024b) 38.5 21.1 39.0 21.4 37.1 19.8
Uni-NaVid* (Zhang et al., 2024a) 41.3 21.1 43.9 21.8 39.5 19.8
MTU3D* (Zhu et al., 2025) 55.0 23.6 45.0 14.7 40.8 12.1
@@
Metrics We evaluate navigation performance using success rate (SR), oracle success rate (OS), success weighted by path length (SPL), and navigation error (NE). Our evaluation protocol is consistent with prior work (Zhang et al., 2024a; Zhu et al., 2025) and follows standard practice.
@@
Instruct goal As shown in Table 1, we compare against all relevant models on the R2R-CE and RxR-CE benchmarks, including both discrete action prediction and waypoint-based continuous prediction, OmniNav achieves state-of-the-art success rates on both benchmarks. Compared with the current best model, OmniNav improves success rate by 4.4% on R2R-CE and 4.3% on RxR-CE.
Object goal To further validate open vocabulary generalization, we also compare OmniNav with prior methods on the HM3D-OVON benchmark. As shown in Table 2, OmniNav again achieves the best performance, with the frontier-based slow system design yielding better results.
@@
Point goal We compare point-goal performance on the CityWalker benchmark, a state-of-the-art point-goal method that supports outdoor navigation scenarios. CityWalker adopts MAOE (Mean Average Orientation Error) as an open-set evaluation metric. On this open-set metric, our approach still outperforms the benchmark method (OmniNav:11.53% vs CityWalker:15.23%)
@@
Ablation Study The independent and synergistic contributions of the four key components policies are evaluated as show in Table.1 and Table. 3). 1) Fast system: autoregressively action chunks generation vs. continuous waypoints generation by flow-matching policy. On R2R-CE, RxR-CE, and OVON benchmarks, the degradation are substantial for action chunks. Their semantic tokens (e.g., left, right) align more easily with language, making them suitable for the first-stage training. However, because action chunks are coarse-grained motion control, continuous waypoints are better suited for fine-grained control. 2) Slow system (planning with frontier and long-term visual memory). We primarily compare with/without the slow system on long-horizon exploration in OVON and find the largest improvements here. Once previously explored areas are recorded, the agent reduces redundant exploration and improves efficiency. Moreover, decomposing active exploration into subgoals (e.g., “go to the bedroom first”) and letting the fast system quickly approach each subgoal forms a hierarchical ”plan–execute” loop, which better matches human reasoning and behavior in unfamiliar environments. 3) General data (general MLLM and referring/grounding). With the slow system enabled, adding the general-purpose datasets yields further stable gains. 4) CoT (explicit chain-of-thought outputs). Using CoT makes the basis for subgoal selection in the slow system transparent, enabling process-level self-check and correction. It reduces cumulative errors in long chains and complex semantic tasks, producing stable improvements. When all four are enabled, performance is best: the slow system provides semantically plausible long-horizon subgoals; the policy head executes with high precision and low latency; general data injects commonsense and language–vision alignment; and CoT provides auditable processes and self-correction.
@@
Real-world deployment In real-robot deployment see Fig. 4, we deploy the fast system component of the model architecture—comprising the VLM and the policy head on a cloud server with an RTX 3090 GPU. The history buffer holds up to 20 front-view frames with at a resolution of 120×106, while the current input is tri-view at 480×426. The system runs at over 5 Hz. For all tasks, it outputs waypoints, which are then fed into the onboard speed control module, which, based on the current speed and maximum acceleration, generates a set of candidate speeds. It then selects the optimal speed to approach the waypoint (selection criteria: minimize speed changes to maintain high speed as much as possible, and be closest to the target waypoint)

### v3

@@ 4 DATA and Training
Data in the embodied domain is typically organized as a data pyramid (bjorck2025gr00t), with internet data and human video data at the bottom, simulation or synthetic data in the middle, and real-robot data at the top. Our dataset follows this rich composition as well, including general web data, simulation data, and a very small amount of real-robot data detailed in Fig. 3
@@ 4.1 General Dataset
It is found in our experiment that models can learn the navigation paradigm relatively easily, whereas general-purpose capabilities remain challenging. To strengthen these abilities, we extended the training data with broad-coverage general-purpose datasets. General-purpose datasets (general QA, image captioning, OCR, chart understanding, coding, and math) complement Vision-and-Language Navigation (VLN) by supplying foundational skills in language understanding, visual semantics, OCR recognition, structured reasoning, and algorithmic planning. These capabilities improve instruction comprehension, path planning, and overall robustness. They also introduce commonsense and functional priors, such as “bath towels are commonly found in bathrooms”. We draw on the open-source MAmmoTH-VL corpus (guo2024mammoth), and subsampled examples to expand our general capabilities; the composition ratios are shown in Fig. 3.
We further incorporated grounding and referring data to more reliably ground linguistic targets and relations to concrete pixels, instances, and locations in the scene, thereby yielding policies that are more interpretable, robust, and generalizable. This includes fine-grained language-to-target mapping—for example, precisely localizing language like “red sofa,” “door with a handle,” or “the second chair” to the correct image regions to support instance-level disambiguation. It also encompasses spatial and relational understanding, such as learning to ground spatial prepositions and relations (“to the left/in front of/at the corner”), and extracting actionable turning and stopping cues from relational descriptions like “the door next to the painting” or “after passing the hallway, turn right at the first intersection.” The referring and grounding data are sourced from RefCOCO series (chen2025revisiting) and Objects365 (shao2019objects365).
@@ 4.2 Multi-Task Navigation Dataset
Point-goal data. We use the open-source Citywalker corpus (liu2025citywalker), which consists of first-person city-walking videos collected from public YouTube channels. Each trajectory comprises a single forward-view sequence. Long videos are segmented into 2-minute clips, then the camera pose for every frame are recovered by DPVO (Deep Patch Visual Odometry) (teed2023deep).
Instruct-goal data. Under Habitat’s VLN-CE (continuous environment) setting and Matterport3D scens (chang2017matterport3d), based on the public instruction–path pairs from R2R (anderson2018vision) and RxR (ku2020room). For each trajectory, we store a panoramic sequence composed of three-views (front, left, right), the action and waypoint sequences, and the natural-language instruction.
Object-goal data. Specifically, open-vocabulary object navigation (OVON) data (yokoyama2024hm3d). In Habitat-Matterport3D (HM3D) (ramakrishnan2021habitat) scenes, we randomly sample (start pose, target object category) pairs. The built-in shortest-path navigation algorithm is used to move the agent to the vicinity of the target. The same content are recorded as the instruct-goal data for each trajectory.
Object-goal (with frontier-based exploration) data. At each step the agent updates its occupancy map and visible region to identify all current frontiers. A policy then selects one frontier—favoring the shortest path while introducing limited randomness—as the current subgoal. Once the target object is found, the exploration episode terminates successfully, yielding a trajectory (zhu2025move). Each trajectory record includes a single forward-view frame sequence, a unexplored frontiers sequence and a natural-language instruction.
Embodied QA. ScanQA (azuma2022scanqa) focuses on real-world indoor 3D scene understanding, with QA pairs centered on object locations, attributes, and spatial relations. R2R-EnvDrop (tan2019learning) addresses continuous visual navigation: scenes are from Matterport3D, trajectories are from R2R-CE, and the navigation instruction–trajectory pairs are recast into a QA format to strengthen alignment between linguistic expressions and visual observations.
Navigation data Process.Each action step corresponds to a 3D continuous pose, described jointly by position and orientation. Taking the agent’s current front-view coordinate frame as the origin, all other trajectory points are transformed into this local frame. Then, we project each 3D pose onto the ground plane, keeping only the planar coordinates $(x,y)$ and the heading angle $(\theta)$, and we additionally attach an “arrival" flag to each step. In this way, each trajectory point is ultimately represented as a quadruple $(x,y,\theta,arrive)$, which serves as the primary target for waypoint optimization. Under this abstraction, since the underlying simulator ensure that any valid 2D path passing through a stair region can be realized as a 3D trajectory across floors, this representation naturally supports multi-floor navigation in HM3D: moving from one floor to another is encoded as following a sequence of 2D poses from the stair entrance to the stair exit, while the vertical motion is handled implicitly by the simulator. At the same time, the agent always receives 3D visual observations with full geometric information, so the stair structure and floor changes are reflected in the visual features.
@@ 4.3 Discrete and Continuous Joint Training
A joint training scheme is particularly critical for Vision-and-Language Navigation (VLN), which requires strong general knowledge and semantic understanding, leading to substantial improvements in success rates in open environments. Stage 1 is trained with 96 NVIDIA H20 GPUs for 120 hours, and Stage 2 is trained with 64 NVIDIA H20 GPUs for 48 hours with lower learning rate.
@@ 5 Experiments
HPN+DN* (krantz2021waypoint) 6.31 40.0 36.0 34.0 - - -
CMA* (hong2022bridging) 6.20 52.0 41.0 36.0 8.76 26.5 22.1
Sim2Sim* (krantz2022sim) 6.07 52.0 43.0 36.0 - - -
GridMM* (wang2023gridmm) 5.11 61.0 49.0 41.0 - - -
DreamWalker* (wang2023dreamwalker) 5.53 59.0 49.0 44.0 - - -
Reborn* (an20221st) 5.40 57.0 50.0 46.0 5.98 48.6 42.0
ETPNav* (an2024etpnav) 4.71 65.0 57.0 49.0 5.64 54.7 44.8
HNR* (wang2024lookahead) 4.42 67.0 61.0 51.0 5.50 56.3 46.7
AG-CMTP (chen2021topological) 7.90 39.0 23.0 19.0 - - -
R2R-CMTP (chen2021topological) 7.90 38.0 26.0 22.0 - - -
InstructNav (long2024instructnav) 6.89 - 31.0 24.0 - - -
LAW (raychaudhuri2021language) 6.83 44.0 35.0 31.0 10.90 8.0 8.0
CM2 (georgakis2022cross) 7.02 41.0 34.0 27.0 - - -
WS-MGMap (chen2022weakly) 6.28 47.0 38.0 34.0 - - -
AO-Planner (chen2025affordances) 5.55 59.0 47.0 33.0 7.06 43.3 30.5
Seq2Seq (krantz2020beyond) 7.77 37.0 25.0 22.0 12.10 13.9 11.9
CMA (krantz2020beyond) 7.37 40.0 32.0 30.0 - - -
NaVid (zhang2024navid) 5.47 49.0 37.0 35.0 - - -
Uni-NaVid (zhang2024uni) 5.58 53.5 47.0 42.7 6.24 48.7 40.9
NaVILA (cheng2024navila) 5.22 62.5 54.0 49.0 6.77 49.3 44.0
StreamVLN (wei2025streamvln) 4.98 64.2 56.9 51.9 6.22 52.9 46.0
CorrectNav (yu2025correctnav) 4.24 67.5 65.1 62.3 4.09 69.3 63.3
@@
Method Observation Val-Seen Val-Seen-Synonyms Val-Unseen
@@
BC 11.1 4.5 9.9 3.8 5.4 1.9
DAgger 11.1 4.5 9.9 3.8 5.4 1.9
RL 18.1 9.4 15.0 7.4 10.2 4.7
DAgRL 41.3 21.2 29.4 14.4 18.3 7.9
BCRL 39.2 18.7 27.8 11.7 18.6 7.5
VLFM* (yokoyama2024vlfm) 35.2 18.6 32.4 17.3 35.2 19.6
DAgRL+OD (yokoyama2024hm3d) 38.5 21.1 39.0 21.4 37.1 19.8
Uni-NaVid* (zhang2024uni) 41.3 21.1 43.9 21.8 39.5 19.8
MTU3D* (zhu2025move) 55.0 23.6 45.0 14.7 40.8 12.1
@@
Metrics We evaluate navigation performance using success rate (SR), oracle success rate (OS), success weighted by path length (SPL), and navigation error (NE). Our evaluation protocol is consistent with prior work (zhang2024uni; zhu2025move) and follows standard practice.
@@
Instruct goal As shown in Table 1, On the R2R-CE and RxR-CE benchmarks, we compare our model against all relevant competitors, including both discrete and continuous prediction methods. Notably, using only its fast system and pure RGB inputs, OmniNav achieves state-of-the-art success rates on both benchmarks. It surpasses the previous leading model by improving the success rate by 4.4% on R2R-CE and 4.3% on RxR-CE.
Object goal To further validate open vocabulary generalization, we also compare OmniNav with prior methods on the HM3D-OVON benchmark. As shown in Table 2, under purely visual inputs, OmniNav already surpasses the best existing approach by 2.7%. However, given that the OVON task demands long-horizon and global planning, the limitations of a purely reactive fast system—such as getting trapped in local loops and exhibiting poor map coverage—become particularly pronounced. Therefore, we incorporate the slow system integrated with frontier-based reasoning, which necessitates the use of depth and odometry information to build and maintain an occupation map. This augmentation equips the agent with global spatial awareness and the capacity for proactive exploration, ultimately leading to superior overall performance(exceeding the strongest prior method by 18.4%).
@@
Point goal We compare point-goal performance on the CityWalker benchmark, a state-of-the-art point-goal method that supports outdoor navigation scenarios. CityWalker adopts MAOE (Mean Average Orientation Error) as an open-set evaluation metric. On this open-set metric, our approach still outperforms the benchmark method (OmniNav:11.53% vs CityWalker:15.23%).
@@
Ablation Study The independent and synergistic contributions of the four key components policies are evaluated as show in Table.1 and Table.3. 1) Fast system: autoregressively action chunks generation vs. continuous waypoints generation by flow-matching policy. On R2R-CE, RxR-CE, and OVON benchmarks, the degradation is substantial for action chunks. Their semantic tokens (e.g., left, right) align more easily with language, making them suitable for the first-stage training. However, because action chunks are coarse-grained motion control, continuous waypoints are better suited for fine-grained control. 2) Slow system (planning with frontier and long-term visual memory). We primarily compare with/without the slow system on long-horizon exploration in OVON and find the largest improvements here. Once previously explored areas are recorded, the agent reduces redundant exploration and improves efficiency. Moreover, decomposing active exploration into subgoals (e.g., “go to the bedroom first”) and letting the fast system quickly approach each subgoal forms a hierarchical "plan–execute" loop, which better matches human reasoning and behavior in unfamiliar environments. 3) General data (general MLLM and referring/grounding). With the slow system enabled, adding the general-purpose datasets yields further stable gains. 4) CoT (explicit chain-of-thought outputs). Using CoT makes the basis for subgoal selection in the slow system transparent, enabling process-level self-check and correction. It reduces cumulative errors in long chains and complex semantic tasks, producing stable improvements. When all four are enabled, performance is best: the slow system provides semantically plausible long-horizon subgoals; the policy head executes with high precision and low latency; general data injects commonsense and language–vision alignment; and CoT provides auditable processes and self-correction.
@@
Real-world deployment In real-robot deployment see Fig. 4, we deploy the fast system component of the model architecture—comprising the VLM and the policy head on a cloud server with an RTX 3090 GPU. The history buffer holds up to 20 front-view frames with at a resolution of 120×106, while the current input is tri-view at 480×426. The system runs at over 5 Hz. For all tasks, it outputs waypoints, which are then fed into the onboard speed control module, which, based on the current speed and maximum acceleration, generates a set of candidate speeds. It then selects the optimal speed to approach the waypoint (selection criteria: minimize speed changes to maintain high speed as much as possible, and be closest to the target waypoint).
Deploying the full slow system in real-world settings requires additional engineering, such as robust real-time integration with LiDAR/depth estimation. This paper mainly focuses on validating the effectiveness of the dual-system collaboration framework in terms of navigation performance and behavior. A full physical deployment of the complete slow system, and systematic optimization of its performance under real-world constraints (including detailed latency–frequency trade-off analysis), constitutes an important avenue for future research.
@@ 6 CONCLUSION & FUTURE WORKS
Appendix A Appendix
A.1 Analysis of Training Data and Model scale
Analysis of Training Data The three additional data components shown in Figure 3 are ablated individually. A qualitative analysis of specific examples reveals the following: 1) Removing Embodied Q&A or Grounding/referring noticeably degrades performance on small background objects (e.g., “picture", “flowerpot"), suggesting these data improve recognition of small objects. 2) Removing General MLLM data leads to failures on irregular objects (e.g., “handrail", “stair"), implying general vision–language data helps with such targets. These findings motivate constructing specific datasets to address failure modes. For example, the model still struggles with “carpet" and “clothes" where complex folds and textures are involved; we therefore can curate dedicated VQA-style data emphasizing these patterns to strengthen fine-grained visual reasoning on texture-heavy objects.
Table 4: Ablation Study of Data on HM3D-OVON Val-Unseen
Embodied Q&A Data Grounding and Referring Data General MLLM Data Ovon-Unseen
55.9
56.5
56.7
57.0
57.7
Analysis of Model Scale The influence of model size (Embodied Q&A, Grounding and Referring, and General MLLM data) is also analyzed. When including these additional data, the 3B and 7B models exhibit nearly identical navigation performance, indicating that once such data are incorporated, simply scaling up the model size brings little further improvement.
In contrast, in the absence of this additional data, a noticeable performance gap emerges, with the 7B model outperforming the 3B model. Our analysis suggests two key factors: 1) the performance of the 3B model appears to be constrained by data sufficiency rather than its inherent capacity. When provided with diverse and abundant data, its performance becomes comparable to that of the 7B model, indicating that model size itself is not the primary bottleneck in this data-rich scenario. 2) based on an analysis of failure cases, we find that further performance gains are limited by the intrinsic difficulty of the task: regardless of model size, the recognition of complex objects such as clothes and mirrors remains unstable.
Beyond these two ablation study, a more systematic study of scaling laws and optimal training configurations—e.g., how data quality, data composition, and more model size jointly affect performance—would also be highly valuable. Due to computational limits we have not yet conducted such a systems-level exploration, and we view this as an important direction for future work.
Table 5: Ablation Study of Model Size on HM3D-OVON Val-Unseen
Model Additional Data Ovon-Unseen
3B 55.9
7B 57.2
3B 57.7
7B 57.9